0 1 pts Question 9 The distribution is normal, the mean is 60 and the standard deviation is 5. What percent of scores lie above 55? -0.8413 0.1587 15.87 84.13
Answer #1

\mu = 60

\sigma= 5

To find P(X>55):

Z= (55 - 60)/5 = - 1

Table of Area Under Standard Normal Curv gives area = 0.3413

So,

P(X>55) = 0.5 + 0.3413 = 0.8413 = 84.13 %

So,

Correct option:

'84.13